OpenStudy (anonymous):

I need help! I suck at math, how do you find the sin of 11pi/3 without a calculator??

OpenStudy (anonymous):

First remember to get it below \(2\pi\).

OpenStudy (anonymous):

\[ 11\pi/3 = (6\pi+5\pi)/3 = 2\pi+5\pi/3 \]So the angle is just \(5\pi/3\)

OpenStudy (anonymous):

uhh okay i kinda get that! but what about the rest??

OpenStudy (anonymous):

Since it's greater than \(\pi\), we can change it to a negative angle: \[ 5\pi/3 -2\pi = -\pi/3 \]

OpenStudy (anonymous):

how did you just take away the five like that?

OpenStudy (whpalmer4):

\[2\pi = \frac{6\pi}{3}\]

OpenStudy (anonymous):

You can add/subtract \(2\pi\) from an angle and the result will be the same because it is the period of the \(\sin\) function.

OpenStudy (anonymous):

AND IT ALL MAKES SENSE NOW!!!!

OpenStudy (anonymous):

Also remember that \(\sin\) is odd \[ \sin(-\pi/3) = -\sin(\pi/3) \]

OpenStudy (anonymous):

okay, so i have that, but i need like a fraction, and it doesn't seem to be coming out evenly...

OpenStudy (anonymous):

how did you know the bottom and side numbers? (sorry if i'm bothering you) i'm trying to figure out my math homework lol

OpenStudy (anonymous):

I made a mistake.

OpenStudy (anonymous):

and according to my math homework website the correct answer was negative radical 3 over 2...

OpenStudy (anonymous):

I see that now haha

OpenStudy (anonymous):

Well, start with an equilateral triangle. All sides are length 1, all angles add up to \(\pi\), so they are \(\pi/3\).

OpenStudy (anonymous):

Then we cut it in half, to form a right triangle.